Find the x and y intercepts and graph the equation plotting the intercepts
-3x-7y=21

Answer by ewatrrr(24785)


Hi,
-3x-7y=21 OR y = (-3/7)x -3
x-intercept (y = 0)
-3x = 21
x = -7 Pt(-7,0)
y-intercept (x=0)
-7y = 21
y = -3 Pt(0,-3)
graphing